OpenStudy (anonymous):

What is the exact diameter of a circular rug with a circumference of 21 feet?

OpenStudy (jonathanvaldes98):

circ = pi *dia 72=22/7* dia dia = 72*7/22 dia = 22.91 ft.

OpenStudy (anonymous):

ty

OpenStudy (jonathanvaldes98):